# Self-Contrast: Better Reflection Through Inconsistent Solving Perspectives

### v2

@@ Abstract
The reflection capacity of Large Language Model (LLM) has garnered extensive attention. A post-hoc prompting strategy, e.g., reflexion and self-refine, refines LLM’s response based on self-evaluated or external feedback. However, recent research indicates without external feedback, LLM’s intrinsic reflection is unstable. Our investigation unveils that the key bottleneck is the quality of the self-evaluated feedback. We find LLMs often exhibit overconfidence or high randomness when self-evaluate, offering stubborn or inconsistent feedback, which causes poor reflection. To remedy this, we advocate Self-Contrast: It adaptively explores diverse solving perspectives tailored to the request, contrasts the differences, and summarizes these discrepancies into a checklist which could be used to re-examine and eliminate discrepancies. Our method endows LLM with diverse perspectives to alleviate stubborn biases. Moreover, their discrepancies indicate potential errors or inherent uncertainties that LLM often overlooks. Reflecting upon these can catalyze more accurate and stable reflection. Experiments conducted on a series of reasoning and translation tasks with different LLMs serve to underscore the effectiveness and generality of our strategy.
@@
Wenqi Zhang1, Yongliang Shen1, Linjuan Wu1 Qiuying Peng2, Jun Wang2, Yueting Zhuang1, Weiming Lu1 1College of Computer Science and Technology, Zhejiang University 2OPPO Research Institute, China {zhangwenqi, luwm}@zju.edu.cn
@@ 1 Introduction
  • We advocate Self-Contrast: LLMs create multiple solving perspectives for diverse results, mitigating overconfident biases of a singular prompt. Then drawing inspiration from contrasting different perspectives, LLM summarizes more accurate checking instructions to resolve discrepancies and enhance reflection.
@@ 2.1 Performance Pre- and Post-Reflection
$t(\Delta\!>\!0)$uparrow -1.94 -0.2 -0.7
Table 1: We calculate the average accuracy of the ten experiments for pre- and post-reflection: Pre Acc. $\Rightarrow$ Post Acc. We also report the significance level of the accuracy change for ten trials, where $\Delta\!=\!Post-Pre$.
@@ 2.2 Feedback Analysis
To investigate the reasons behind the failure of reflection, we further analyze the feedback generated during the self-evaluate process. We classify all samples in GSM8K into four categories based on their correctness of the pre- and post-reflection: 1) Invalid Reflection ($\Rightarrow$) means the results before and after reflection are both incorrect. 2) Valid Reflection ($\Rightarrow$) means a wrong solution is revised to correct through reflection. 3) Toxic Reflection ($\Rightarrow$) represents that an originally correct response is changed to incorrect after reflection. 4) Others counts the number of correct $\Rightarrow$ correct.
@@
Often Provide Overconfident or Inconsistent Feedback. We examine whether LLMs could generate feedback accurately and consistently. For each sample, we instruct LLM to evaluate its initial response multiple times and record multiple feedbacks. We manually assess the consistency and correctness of these feedbacks and then summarize each sample into 4 cases: I. Accurately identifies errors: In multiple repeated evaluations, LLM identifies errors and provides accurate and consistent feedback. II. Stubbornly offers erroneous feedback: The majority of evaluations provide incorrect feedback with specific errors. III. Can not output consistent feedback: Unable to assess consistently, as most feedback is different and quite random for a same initial response. V. Overconfidence, no revision required: LLM is overconfident and believes no revision is necessary.
@@ 2.3 From Self-Evaluate to Self-Contrast
As a remedy, we propose a contrastive strategy for reflection. Instead of directly evaluating a response, which can be challenging and inconsistent, we instruct LLM to initially contrast the differences between various solutions, and identify their discrepancies and the reasons behind them. As shown in Figure 1 (bottom), we sample Top-2 responses from LLM and then prompt LLM to contrast their differences in detail, rethink the reasons that caused the discrepancies, and summarize the checklist for re-examining and resolving the discrepancy. As shown in Table 3, we compare three scenarios: self-evaluate w/ top-1 response, self-evaluate w/ top-2 responses, and self-contrast w/ top-2. Our new strategy achieves a modest improvement over standard reflection using self-evaluate. Notably, it significantly enhances the significance levels (t-test: -0.43 to 1.43), suggesting it can greatly mitigate the uncertainty associated with self-evaluate process.
@@ 3 Self-Contrast
Prior sections illustrate the challenges LLMs encounter in accurately evaluating previous solutions, often resulting in overconfident or inconsistent feedback. Concurrently, we observe that leveraging the discrepancies between two different solutions can catalyze a more efficacious reflection, notably reducing the uncertainty during the reflection. Building upon this insight, we propose a more diverse inter-perspective Self-Contrast, facilitating more reliable self-reflection.
@@ 4.1 Settings
LLM Models and Prompts We conduct experiments using the GPT-3.5-Tubor-0613 and GPT-4-0613, alongside the Llama2-Chat model with three parameter scales (7B, 13B, and 70B). To make a fair comparison, we uniformly set the temperature to 0.2 for all experiments. For standard prompts and self-reflection baseline, we evaluate them 10 times using different prompts and average their results under zero-shot scenes. Prompts and other details can be found in Appendices C, D and B.2.
@@ Appendix A Complementary Experiments
A.1 LLM is More Likely to Trust Previous Response
@@
A.2 Self-Evaluate Vs. Self-Contrast
@@
A.3 Ablation Study For Selection Strategy
@@ E.4 Learning Mathematical Reasoning
In recent years, mathematical reasoning has become a significant benchmark (Cobbe et al., 2021; Hendrycks et al., 2021) to evaluate the capabilities of artificial intelligence models. Within the paradigm of supervised learning, a vast amount of research (Xie and Sun, 2019; Patel et al., 2021; Jie et al., 2022; Zhang et al., 2023b) has been dedicated to translating human language into mathematical equations. In the era of LLMs, the advent of Chain-of-Thought and other prompting strategies have notably augmented the reasoning capabilities (Zhu et al., 2023; Yuan et al., 2023b; Frieder et al., 2023; Zhou et al., 2022).
Prompting Method PAL and Program-of-Thoughts Gao et al. (2022); Chen et al. (2022) separate the computation and reasoning process using code as the intermediate process. Mathprompter, Auto-Model (Imani et al., 2023; Zhao et al., 2023) encourage LLMs to generate diverse reasoning paths in different forms simultaneously, including text (CoT), code (PAL), and symbols (Equation) for a higher confidence answer. Automatic-CoT, Complexity-CoT, Synthetic Prompt and Boosted Prompt (Zhang et al., 2022b; Fu et al., 2022; Shao et al., 2023; Pitis et al., 2023) enhance reasoning performance by optimizing the selection of demonstrations within the prompt. Tree-of-thought and Self-Evaluation (Yao et al., 2023; Xie et al., 2023) extend the CoT into a search tree, obtaining more accurate answers through self-evaluation.

### v3

@@ Abstract
The reflection capacity of Large Language Model (LLM) has garnered extensive attention. A post-hoc prompting strategy, e.g., reflexion and self-refine, refines LLM’s response based on self-evaluated or external feedback. However, recent research indicates without external feedback, LLM’s intrinsic reflection is unstable. Our investigation unveils that the key bottleneck is the quality of the self-evaluated feedback. We find LLMs often exhibit overconfidence or high randomness when self-evaluate, offering stubborn or inconsistent feedback, which causes poor reflection. To remedy this, we advocate Self-Contrast: It adaptively explores diverse solving perspectives tailored to the request, contrasts the differences, and summarizes these discrepancies into a checklist which could be used to re-examine and eliminate discrepancies. Our method provides LLM with diverse perspectives to alleviate stubborn biases. Moreover, their discrepancies indicate potential errors or inherent uncertainties that LLM often overlooks. Reflecting upon these can prompt more accurate and stable reflection. Experiments conducted on a series of reasoning and translation tasks with different LLMs serve to underscore the effectiveness and generality of our strategy.
@@
Wenqi Zhang1, Yongliang Shen1, Linjuan Wu1 Qiuying Peng2, Jun Wang2, Yueting Zhuang1, Weiming Lu1† 1College of Computer Science and Technology, Zhejiang University 2OPPO Research Institute, China {zhangwenqi, luwm}@zju.edu.cn
22footnotetext: Corresponding author.
@@ 1 Introduction
  • We advocate Self-Contrast: the LLMs create multiple solving perspectives for diverse results, mitigating overconfident biases of a singular prompt. Then drawing inspiration from contrasting different perspectives, the LLMs summarize more accurate checking instructions to resolve discrepancies and enhance reflection.
@@ 2.1 Performance Pre- and Post-Reflection
P for $\Delta>0$ 0.9578 0.5770 0.7492
Table 1: We calculate the average accuracy of the ten experiments for pre- and post-reflection: Pre Acc. $\Rightarrow$ Post Acc. We also report the accuracy change’s significance level (P-value) for ten trials, where $\Delta\!=\!Post-Pre$. A larger P indicates a less significant improvement.
@@ 2.2 Feedback Analysis
To investigate the reasons behind the failure of reflection, we further analyze the feedback generated during the self-evaluate process. We classify all samples in GSM8K into four categories based on their correctness of the pre- and post-reflection: 1) Invalid Reflection ($\Rightarrow$) means the results before and after reflection are both incorrect. 2) Valid Reflection ($\Rightarrow$) means a wrong solution is revised to correct through reflection. 3) Toxic Reflection ($\Rightarrow$) represents that an originally correct response is changed to incorrect after reflection. 4) Others counts the number of correct $\Rightarrow$ correct. Automatic statistics for the reflection category.
Step 1: We categorize the reflection into the above four categories. This process can be automated for mathematical benchmarks by comparing whether the answers are correct before and after reflection. For the translation task, we leverage GPT-4 along with annotated answers to evaluate the accuracy of translation results before and after reflection. Step 2: We manually assess the quality of the feedback generated in each reflection case (Invalid, Valid, and Toxic). Based on the correctness and consistency of these feedbacks, we categorize them into four cases (inconsistent, overconfident, etc.). The detailed results are as follows:
@@
Often Provide Overconfident or Inconsistent Feedback. We examine whether LLMs could generate feedback accurately and consistently. For each sample, we instruct the LLM to evaluate its initial response multiple times and record multiple feedbacks. We manually assess the consistency and correctness of these feedbacks and then summarize each sample into 4 cases: I. Accurately identifies errors: In multiple repeated evaluations, the LLM identifies errors and provides accurate and consistent feedback. II. Stubbornly offers erroneous feedback: The majority of evaluations provide incorrect feedback with specific errors. III. Can not output consistent feedback: Unable to assess consistently, as most feedback is different and quite random for a same initial response. V. Overconfidence, no revision required: LLM is overconfident and believes no revision is necessary. The detailed evaluation criteria are provided in Section A.1.
@@ 2.3 From Self-Evaluate to Self-Contrast
As a remedy, we propose a contrastive strategy for reflection. Instead of directly evaluating a response, which can be challenging and inconsistent, we instruct the LLM to initially contrast the differences between various solutions, and identify their discrepancies and the reasons behind them. As shown in Figure 1 (bottom), we sample Top-2 responses from LLM and then prompt LLM to contrast their differences in detail, rethink the reasons that caused the discrepancies, and summarize the checklist for re-examining and resolving the discrepancy. As shown in Table 3, we compare three scenarios: self-evaluate w/ top-1 response, self-evaluate w/ top-2 responses, and self-contrast w/ top-2. Our new strategy achieves a modest improvement over standard reflection using self-evaluate. Notably, it significantly enhances the significance levels (p-value: 0.6613 to 0.0933), suggesting it can greatly mitigate the self-evaluation process’s uncertainty.
In this section, we validate the concept of contrastive evaluation. For the next section, we expand this contrastive concept into full-version self-contrast, which involves creating multiple perspectives, and contrasting their differences, summarizing the checklist for deeper reflection.
@@ 3 Self-Contrast
Prior sections illustrate the challenges LLMs encounter in accurately evaluating previous solutions, often resulting in overconfident or inconsistent feedback. Concurrently, we observe that leveraging the discrepancies between two different solutions can inspire a more efficacious reflection, notably reducing the uncertainty during the reflection. Building upon this insight, we propose a more diverse inter-perspective Self-Contrast, facilitating more reliable self-reflection.
@@ 4.1 Settings
LLM Models and Prompts We conduct experiments using the GPT-3.5-Turbo-0613 and GPT-4-0613, alongside the Llama2-Chat model with three parameter scales (7B, 13B, and 70B). To make a fair comparison, we uniformly set the temperature to 0.2 for all experiments. For standard prompts and self-reflection baseline, we evaluate them 10 times using different prompts and average their results under zero-shot scenes. Prompts and other details can be found in Appendices C, D and B.2.
@@ 6.3 Diverse Solving Perspectives Maximize Contrast Effect
7 Discussions
Self-contrast switches the critique objective into a contrastive task. We transform the self-evaluation into a process of comparing differences, explicitly altering the attention distribution of the LLM. The LLM is required only to identify the differences between two solutions, without judging right or wrong. This process is less influenced by the biases inherent in LLMs, as the objective is contrasting rather than evaluation. Besides, in Table 7, LLMs are instructed to contrast two incorrect solutions with different errors which also improves reflection results.
The results in Section 2.3 also precisely verify this conjecture. By simply transforming from direct evaluation to contrastive evaluation, we enhance the effectiveness of reflection (75.8 to 77.5 on GSM8K), with more significant results (0.66 to 0.09). In Tables 4 and 5, our self-contrast approach achieved more significant improvements.
Contrasting results can help LLMs notice overlooked details and biases. After contrasting the differences between the two solutions, we summarize these differences into a checklist, thereby explicitly prompting the LLM to focus on the logical pitfalls and other issues underlying these differences. This allows LLMs to engage in reflection more clearly and purposefully.
As shown in Figure 2, LLM generates different translations for the user’s request: "This plan was shot to death", and "This plan was axed". The former is a rigid translation that fails to grasp the metaphor embedded in the military term. After contrasting two different translations, LLMs believe they should scrutinize the source sentence for metaphors and ensure the translation aligns with the conventions of English expression.
@@ 8 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 62376245), the Key Research and Development Program of Zhejiang Province, China (No. 2024C01034), the Fundamental Research Funds for the Central Universities, and National Key Research and Development Project of China (No. 2018AAA0101900).
@@ Appendix A Complementary Experiments
A.1 Detail For Manual Feedback Evaluation
We provide the details of manual evaluation in Table 2. Specifically, we categorize reflection results into four categories: Invalid (wrong -> wrong), Valid (wrong -> right), Toxic (right -> wrong), and Other (right -> right) based on their answer. Subsequently, we manually assess the quality of feedback within each reflection category. For instance, for a Toxic reflection case, we devise ten self-evaluation prompts, prompting LLMs to conduct a self-evaluation on their initial response, generating ten feedbacks. These feedbacks are manually checked for correctness and consistency. The criteria for classification are as follows:
  • If more than seven feedbacks accurately identify the errors in the initial response, we classify it into the first category: I. Accurately Identifies Errors. Please note that the feedback only needs to accurately identify the errors without necessarily correcting them.
  • If more than seven feedbacks indicate that the initial response has no errors, we categorize it into the fourth category: IV. Overconfidence - No Revision Required.
  • If, among the ten feedbacks, there are more than three different opinions, e.g., the first feedback suggesting there are no errors, another identifying error-1, and yet another pointing out error-2…. For this scenario, we classify it into the third category: III. Cannot Output Consistent Feedback. In this scenario, self-evaluation exhibits significant randomness.
  • The remaining cases are classified into the second category: II. Stubbornly Offers Erroneous Feedback.
The entire human evaluation process is conducted by two senior PhD students. One is responsible for categorization, while the other verifies the categorization again.
A.2 LLM is More Likely to Trust Previous Response
@@
A.3 Self-Evaluate Vs. Self-Contrast
@@
A.4 Ablation Study For Selection Strategy
@@ E.4 Learning Mathematical Reasoning
Mathematical reasoning is the key to achieving embodied intelligence (Zhang et al., 2021, 2022c). In recent years, mathematical reasoning has become a significant benchmark (Cobbe et al., 2021; Hendrycks et al., 2021) to evaluate the capabilities of artificial intelligence models. Within the paradigm of supervised learning, a vast amount of research (Xie and Sun, 2019; Patel et al., 2021; Jie et al., 2022; Zhang et al., 2022b, 2023b) has been dedicated to translating human language into mathematical equations. In the era of LLMs, the advent of Chain-of-Thought and other prompting strategies have notably augmented the reasoning capabilities (Zhu et al., 2023; Yuan et al., 2023b; Frieder et al., 2023; Zhou et al., 2022).
Prompting Method PAL and Program-of-Thoughts Gao et al. (2022); Chen et al. (2022) separate the computation and reasoning process using code as the intermediate process. Mathprompter, Auto-Model (Imani et al., 2023; Zhao et al., 2023) encourage LLMs to generate diverse reasoning paths in different forms simultaneously, including text (CoT), code (PAL), and symbols (Equation) for a higher confidence answer. Automatic-CoT, Complexity-CoT, Synthetic Prompt and Boosted Prompt (Zhang et al., 2022d; Fu et al., 2022; Shao et al., 2023; Pitis et al., 2023) enhance reasoning performance by optimizing the selection of demonstrations within the prompt. Tree-of-thought and Self-Evaluation (Yao et al., 2023; Xie et al., 2023) extend the CoT into a search tree, obtaining more accurate answers through self-evaluation.